A woman's age is three times her daughters age. In ten years, she will only be twice her daughters age. What are their present ages?
Answer by mananth(16946)

daughter's age now be x
mother's age now = 3x
..
After 10 years
daughter's age x+10
mother's age = 3x+10
..
mother's age = 2 times daughter's age
3x+10= 2(x+10)
3x+10=2x+20
add -10
3x+10-10=2x+20-10
3x=2x+10
add -2x
3x-2x=2x-2x+10
x=10 daughter's age now.
mother's age = 3x = 3*10=30 years